# The cohomological Kudla conjecture for unitary Shimura varieties

### v1

@@ 1.4. Organization of the paper
In Section 2, we give the necessary background on vector-valued Hermitian modular forms and introduce quasi-Hermitian modular forms which come from weighted theta series. We show that the space of weighting polynomials carries an $\mathfrak{sl}_{2}$-action and explain a general method for correcting quasi-modular theta series.
In Section 3, we review the theory of unitary Shimura varieties associated to unitary groups of signature $(n+1,1)$ and their toroidal compactifications. In Section 3.3, we analyze the structure of the boundary, introduce the special cycles in the boundary, and relate the normal bundle of the boundary to the class of the polarization induced by the lattice $(L,h)$. In Section 3.4, we prove the key Splitting Lemma for the rational homology of a toroidal compactification.
In Section 4, we construct harmonic representatives of the special cycles in the boundary that are expressed using homogeneous polynomials valued in the algebra differential forms, and we use Shimura’s modularity criterion to construct the non-holomorphic completions of the generating series of special cycles. We use the $\mathfrak{sl}_{2}$-action to construct the modular corrections.
Combining these tools, the main results of the paper are proved in Section 5. The arguments for the correction and the completion are brief and entirely parallel, thanks to the setup of the preceding sections.
@@ 1.5. Acknowledgments
We would like to thank Benjamin Howard and Wei Zhang for useful discussions, and Philip for helpful discussions and comments. F.G. was supported by NSF grant DMS-2302548, and S.T. was supported by NSF grant DMS-2302388 and DMS-2503815.
@@ 2.1. Hermitian Modular forms
Let $k$ be an imaginary quadratic field with ring of integers ${\mathcal{O}}_{k}$ and different ideal $\mathfrak{D}_{k}$. Let $g\geq 1$ and for a matrix $N\in\mathrm{M}_{g}(\mathbb{C})$, let $N^{*}=\overline{N}^{t}$ denote the conjugate-transpose of $N$.
@@
$\mathrm{U}(g,g)(R)=\{N\in\mathrm{M}_{2g}(R\otimes_{\mathbb{Q}}k)|N^{*}JN=J\}\leavevmode\nobreak\ ,$
@@
$J=\begin{pmatrix}0&\mathrm{I}_{g}\\ -\mathrm{I}_{g}&0\\ \end{pmatrix}\leavevmode\nobreak\ .$
@@
$\mathcal{H}_{g}=\{\tau\in\mathrm{M}_{g}(\mathbb{C})\,|\,Y=\frac{\tau-\tau^{*}}{2i}>0\}\leavevmode\nobreak\ .$
@@
$\tau$ $\mapsto T\cdot\tau=(A\tau+B)(C\tau+D)^{-1}\leavevmode\nobreak\ .$
@@
for every element $M=\begin{pmatrix}A&B\\ C&D\\ \end{pmatrix}\in\mathrm{U}(g,g)(\mathbb{Z})\leavevmode\nobreak\ .$
@@ A Hermitian modular form $f$ admits a Fourier expansion:
$f(\tau)=\sum_{N\in\mathrm{Herm}_{g}(k)}c(N)q^{N},\quad q^{N}=e^{2\pi i\mathrm{Tr}(\tau N)}\leavevmode\nobreak\ ,$
where $\mathrm{Herm}_{g}(k)$ denotes the group of $g\times g$ Hermitian matrices. Since we will be primarily interested in modular forms arising as theta series for possibly non-unimodular Hermitian lattices, we must consider level subgroups of $\mathrm{U}(g,g)(\mathbb{Z})$, or alternatively vector-valued modular forms which we now introduce; see [HRZ24, Section 2.2] for more details.
Let $(L,h)$ be a Hermitian lattice over $\mathcal{O}_{k}$ with a Hermitian form of signature $(p,q)$. Then $(L,\mathrm{Tr}_{k/\mathbb{Q}}(h))$ is naturally a $\mathbb{Z}$-lattice with a quadratic form of signature $(2p,2q)$. The dual lattice $L^{\vee}$ of the $\mathbb{Z}$-lattice of $L$ is isomorphic to its dual as $\mathcal{O}_{k}$lattice and admits the following equivalent descriptions, see [Zem23, Corollary 1.12]:
$L^{\vee}$ =$=\{x\in L_{\mathbb{Q}}\,|\,\forall y\in L,\quad h(x,y)\in\mathfrak{D}^{-1}_{k}\}$
=$=\{x\in L_{\mathbb{Q}}\,|\,\forall y\in L,\quad\mathrm{Tr}_{k/\mathbb{Q}}\left(h(x,y)\right)\in\mathbb{Z}\}\leavevmode\nobreak\ .$
@@
$m(A)$ $:=\begin{pmatrix}A&0\\ 0&(A^{*})^{-1}\end{pmatrix},\quad A\in\mathrm{GL}_{g}({\mathcal{O}}_{k})\leavevmode\nobreak\ ,$
$n(B)$ $:=\begin{pmatrix}\mathrm{I}_{g}&B\\ 0&\mathrm{I}_{g}\end{pmatrix},\quad B\in\mathrm{Herm}_{g}({\mathcal{O}}_{k})\leavevmode\nobreak\ ,$
$w_{g}$ $:=\begin{pmatrix}0&\mathrm{-}I_{g}\\ \mathrm{I}_{g}&0\end{pmatrix}\leavevmode\nobreak\ .$
@@
$\rho_{L,g}(m(A))\mathfrak{e}_{\underline{\nu}}$ $=\det(A)^{-p-q}\mathfrak{e}_{\underline{\nu}A^{-1}}\leavevmode\nobreak\ ,$
@@
for every element $T=\begin{pmatrix}A&B\\ C&D\\ \end{pmatrix}\in\mathrm{U}(g,g)(\mathbb{Z})\leavevmode\nobreak\ .$
@@ 2.2. Homogeneous polynomials and modularity
$P(UA)=|\det(A)|^{2}P(U),\quad\textrm{for all}\quad U\in\mathrm{M}_{n\times g}(\mathbb{C}),\,A\in\mathrm{M}_{g}(\mathbb{C})\leavevmode\nobreak\ .$
@@
$P(\lambda_{1},\ldots,\lambda_{g})=Q(\lambda_{1},\ldots,\lambda_{g},\overline{\lambda}_{1},\ldots,\overline{\lambda}_{g})\leavevmode\nobreak\ .$
@@
$\lambda_{j}\cdot\frac{\partial P}{\partial\lambda_{i}}=\delta_{i,j}P\quad,\quad\overline{\lambda}_{j}\cdot\frac{\partial P}{\partial\overline{\lambda}_{i}}=\delta_{i,j}P\leavevmode\nobreak\ .$
@@
where $Q(X_{1},\ldots,X_{g},Y_{1},\ldots,Y_{g})$ satisfies the condition in Equation 2.1. We then apply [Roe21, Proposition 3.4]
@@
$\lambda_{\ell}=\begin{pmatrix}\frac{\partial}{\partial\lambda_{\ell}^{(1)}}\\ \vdots\\ \frac{\partial}{\partial\lambda_{\ell}^{(n)}}\end{pmatrix}\in\mathbb{C}^{n},$
@@
$\frac{\partial}{\partial\lambda_{\ell}}=\begin{pmatrix}\frac{\partial}{\partial\lambda_{\ell}^{(1)}}\\ \vdots\\ \frac{\partial}{\partial\lambda_{\ell}^{(n)}}\end{pmatrix}\leavevmode\nobreak\ .$
@@
    For any $1\leq s,\ell\leq g$, we have: $\Delta_{s}P=\Delta_{\ell}P$ in ${\mathcal{F}}_{n,g-1}$ as functions of $g-1$ vectors.
@@
$P(\lambda_{1},\ldots,\lambda_{g})=Q(\lambda_{1},\ldots,\lambda_{g},\overline{\lambda}_{1},\ldots,\overline{\lambda}_{g})\leavevmode\nobreak\ .$
@@
$\Delta_{g}P(\lambda_{1},\ldots,\lambda_{g})=\sum_{s=1}^{n}\frac{\partial^{2}Q}{\partial X_{g}^{(s)}\partial Y_{g}^{(s)}}(\lambda_{1},\ldots,\lambda_{g},\overline{\lambda}_{1},\ldots,\overline{\lambda}_{g})\leavevmode\nobreak\ ,$
@@
$Q\left[(X_{1},\ldots,X_{g-1})\cdot A,X_{g},(Y_{1},\ldots,Y_{g-1})\cdot B,Y_{g}\right]\\ =Q\left[(X_{1},\ldots,X_{g})\cdot A_{0},(Y_{1},\ldots,Y_{g})\cdot B_{0}\right]\leavevmode\nobreak\ \\ =\det(A_{0})\det(B_{0})Q(X_{1},\ldots,X_{g},Y_{1},\ldots,Y_{g})\\ =\det(A)\det(B)Q(X_{1},\ldots,X_{g},Y_{1},\ldots,Y_{g})\leavevmode\nobreak\ .$
@@
For $(2)$, without loss of generality, we can assume that $s=g$ and $\ell=g-1$. Let $\sigma$ be the matrix of the transposition $(g-1,g)$. Let $u_{1},\ldots,u_{g-1},\lambda_{g}\in\mathbb{C}^{n}$. The invariance property applied to $\sigma$ ensures that:
$P(u_{1},\ldots,u_{g-1},\lambda_{g})=P(u_{1},\ldots,\lambda_{g},u_{g-1})\leavevmode\nobreak\ .$
@@
$\Delta_{g}P(u_{1},\ldots,u_{g-1})=\Delta_{g-1}P(u_{1},\ldots,u_{g-1})\leavevmode\nobreak\ ,$
@@
Lemma 2.6 above ensures that there exists a well-defined lowering operator for $1\leq g\leq n$:
$\Delta_{g}:{\mathcal{F}}_{n,g}\rightarrow{\mathcal{F}}_{n,g-1}\leavevmode\nobreak\ .$
@@
$\ \Delta:{\mathcal{F}}_{n,\bullet}\rightarrow{\mathcal{F}}_{n,\bullet}\leavevmode\nobreak\ ,$
@@
$m_{ij}=-\prescript{t}{}{\lambda_{i}}\cdot\frac{\partial m_{ii}}{\partial\lambda_{j}}=\sum_{s=1}^{n}{\lambda}_{i}^{(s)}\frac{\partial m_{ii}}{\partial\lambda_{j}^{(s)}}\leavevmode\nobreak\ .$
Using the polynomial $Q$ associated to $P$ as in Equation 2.1 and applying the invariance property for permutation matrices, we can write the $m_{ij}$ as follows:
@@
where we introduce the following notation: for $g$ vectors $\mu_{1},\ldots,\mu_{g}\in\mathbb{C}^{n}$: $[\mu]_{i}^{g}$ is the ordered set of $g-1$ vectors $(\mu_{1},\ldots,\widehat{\mu}_{i},\ldots\mu_{g})$ skipping $\mu_{i}$.
@@
$\lambda_{j}\cdot\frac{\partial m_{\ell k}}{\partial\lambda_{i}}=\lambda_{j}\cdot\frac{\partial}{\partial\lambda_{i}}(P(\lambda_{1},\ldots,\widehat{\lambda}_{\ell},\ldots,\lambda_{g}))\leavevmode\nobreak\ ,$
which 00 if $\ell=i$ (and therefore $k\neq j$). If $\ell\neq i$ and $k\neq j$, then we get 00 by Lemma 2.5. Finally, if $\ell=j(=k)$, then we get $-m_{j,i}=-m_{\ell,i}$.
Assume now that $k\neq\ell$, then $m_{\ell,k}=-\lambda_{\ell}\cdot\frac{\partial m_{\ell\ell}}{\partial\lambda_{k}}$. If $j\neq k$, then $\lambda_{j}\cdot\frac{\partial m_{\ell k}}{\partial\lambda_{i}}=0$, as $m_{\ell,k}$ either does not depend on $\lambda_{i}$, or we can apply Lemma 2.5. Therefore, we may assume that $j=k$. By applying Equation 2.1 to permutation matrices, we get:
@@
$\Lambda(P)(\lambda_{1},\ldots,\lambda_{g})=\mathrm{Tr}\left({{}^{t}{U}}\overline{U}M\right)\leavevmode\nobreak\ ,$
@@
It is enough to check that $\Lambda P$ satisfies the conditions of Lemma 2.5. We can expand
$\Lambda(P)=\sum_{1\leq\ell,s\leq g}h(\lambda_{s},\lambda_{\ell})m_{\ell,s}\leavevmode\nobreak\ .$
@@
$\lambda_{j}\cdot\frac{\partial\Lambda(P)}{\partial\lambda_{i}}=\sum_{1\leq\ell,k\leq g}\lambda_{j}\cdot\frac{\partial}{\partial\lambda_{i}}\left(h(\lambda_{k},\lambda_{\ell})\right)m_{\ell,k}+\sum_{1\leq\ell,k\leq g}h(\lambda_{k},\lambda_{\ell})\lambda_{j}\cdot\frac{\partial m_{\ell,k}}{\partial\lambda_{i}}\leavevmode\nobreak\ .$
@@
$\lambda_{j}\cdot\frac{\partial}{\partial\lambda_{i}}(h(\lambda_{k},\lambda_{\ell}))=\delta_{ik}h(\lambda_{j},\lambda_{\ell})\leavevmode\nobreak\ .$
We distinguish two cases: assume first that $j=i$. Then by Lemma 2.7, (1), we have:
$\lambda_{i}\cdot\frac{\partial m_{\ell,k}}{\partial\lambda_{i}}=(1-\delta_{ik})m_{\ell k}\leavevmode\nobreak\ .$
@@
$=\Lambda P\leavevmode\nobreak\ .$
We now treat the case $i\neq j$. By Lemma 2.7 (2), we have: $\lambda_{j}\cdot\frac{\partial m_{\ell,k}}{\partial\lambda_{i}}=0$ if $k\neq j$ and otherwise it is equal to $-m_{\ell,i}$. Therefore, we get:
@@
$=0\leavevmode\nobreak\ .$
We compute similarly $\overline{\lambda}_{i}\cdot\frac{\partial}{\partial\overline{\lambda_{j}}}\Lambda(P)=\delta_{ij}\Lambda P$. By Lemma 2.5, we get the desired result. ∎
@@
$[\Lambda,\Delta]=H,\quad[H,\Lambda]=2\Lambda,\quad[H,\Delta]=-2\Delta\leavevmode\nobreak\ .$
@@
(2.4) $\Lambda\circ\Delta_{g}(P)=\sum_{1\leq k,\ell\leq g}h(\lambda_{k},\lambda_{\ell})m_{\ell,k}(P_{g})\leavevmode\nobreak\ .$
@@
$+\sum_{1\leq k,\ell\leq g+1}\left[{\frac{{}^{t}\partial}{\partial\lambda_{g+1}}}h(\lambda_{k},\lambda_{\ell})\frac{\partial}{\partial\overline{\lambda}_{g+1}}m_{\ell,k}(P)+{\frac{{}^{t}\partial}{\partial\overline{\lambda}_{g+1}}}h(\lambda_{k},\lambda_{\ell})\frac{\partial}{\partial\lambda_{g+1}}m_{\ell,k}(P)\right]\leavevmode\nobreak\ ,$
@@
    $m_{g+1,g+1}(P)=P$, $\Delta_{g+1}(h(\lambda_{k},\lambda_{\ell}))=n$, if $k=\ell=g+1$ and 00 otherwise.
@@
    $\Delta_{g+1}(m_{g+1,k})=\Delta_{g+1}(m_{\ell,g+1})=0$ for all $11\leq\ell,k\leq g+$ 1, as $m_{g+1,k}$ and $m_{\ell,g+1}$ depend linearly on $\lambda_{g+1}$ and $\overline{\lambda}_{g+1}$ respectively.

### v2

@@ 1.4. Organization of the paper
In Section˜2, we give the necessary background on vector-valued Hermitian modular forms and introduce quasi-Hermitian modular forms which come from weighted theta series. We show that the space of weighting polynomials carries an $\mathfrak{sl}_{2}$-action and explain a general method for correcting quasi-modular theta series.
In Section˜3, we review the theory of unitary Shimura varieties associated to unitary groups of signature $(n+1,1)$ and their toroidal compactifications. In Section˜3.3, we analyze the structure of the boundary, introduce the special cycles in the boundary, and relate the normal bundle of the boundary to the class of the polarization induced by the lattice $(L,h)$. In Section˜3.4, we prove the key Splitting Lemma for the rational homology of a toroidal compactification.
In Section˜4, we construct harmonic representatives of the special cycles in the boundary that are expressed using homogeneous polynomials valued in the algebra differential forms, and we use a modularity criterion to construct the non-holomorphic completions of the generating series of special cycles. We use the $\mathfrak{sl}_{2}$-action to construct the modular corrections.
Combining these tools, the main results of the paper are proved in Section˜5. The arguments for the correction and the completion are brief and entirely parallel, thanks to the setup of the preceding sections.
@@ 1.5. Acknowledgments
We thank Philip Engel, Benjamin Howard, and Wei Zhang for useful discussions and helpful comments on a first draft of this paper. F.G. was supported by NSF grant DMS-2302548, and S.T. was supported by NSF grants DMS-2302388 and DMS-2503815.
@@ 2.1. Hermitian Modular forms
Let $k$ be an imaginary quadratic field with ring of integers ${\mathcal{O}}_{k}$ and different ideal $\mathfrak{D}_{k}$. Let $g\geq 1$ and for a matrix $N\in\mathrm{M}_{g}(\mathbb{C})$, let $N^{*}=\overline{N}^{t}$ denote its conjugate-transpose.
@@
$\mathrm{U}(g,g)(R)=\{N\in\mathrm{M}_{2g}(R\otimes_{\mathbb{Q}}k)|N^{*}JN=J\}\penalty 10000\ ,$
@@
$J=\begin{pmatrix}0&\mathrm{I}_{g}\\ -\mathrm{I}_{g}&0\\ \end{pmatrix}\penalty 10000\$.
@@
$\mathcal{H}_{g}=\{\tau\in\mathrm{M}_{g}(\mathbb{C})\,|\,Y=\frac{\tau-\tau^{*}}{2i}>0\}\penalty 10000\ .$
@@
$\tau$ $\mapsto T\cdot\tau=(A\tau+B)(C\tau+D)^{-1}\penalty 10000\ .$
@@
for every element $T=\begin{pmatrix}A&B\\ C&D\\ \end{pmatrix}\in\mathrm{U}(g,g)(\mathbb{Z})\penalty 10000\ .$
@@ A Hermitian modular form $f$ admits a Fourier expansion:
$f(\tau)=\sum_{N\in\mathrm{Herm}_{g}(k)}c(N)q^{N},\quad q^{N}=e^{2\pi i\mathrm{Tr}(\tau N)}\penalty 10000\ ,$
where $\mathrm{Herm}_{g}(k)$ denotes the group of $g\times g$ Hermitian matrices. Since we will be primarily interested in modular forms arising as theta series for possibly non-unimodular Hermitian lattices, we must consider level subgroups of $\mathrm{U}(g,g)(\mathbb{Z})$, or alternatively vector-valued modular forms which we now introduce; see [BRZ24, Section 2.2] for more details.
Let $(L,h)$ be a Hermitian lattice over $\mathcal{O}_{k}$ with a Hermitian form of signature $(p,q)$. Then $(L,\mathrm{Tr}_{k/\mathbb{Q}}(h))$ is naturally a $\mathbb{Z}$-lattice with a quadratic form of signature $(2p,2q)$. The dual lattice $L^{\vee}$ of the $\mathbb{Z}$-lattice of $L$ is isomorphic to its dual as $\mathcal{O}_{k}$ -lattice and admits the following equivalent descriptions, see [Zem23, Corollary 1.12]:
$L^{\vee}$ $=\{x\in L_{\mathbb{Q}}\,|\,\forall y\in L,\quad h(x,y)\in\mathfrak{D}^{-1}_{k}\}$
$=\{x\in L_{\mathbb{Q}}\,|\,\forall y\in L,\quad\mathrm{Tr}_{k/\mathbb{Q}}\left(h(x,y)\right)\in\mathbb{Z}\}\penalty 10000\ .$
@@
$m(A)$ $:=\begin{pmatrix}A&0\\ 0&(A^{*})^{-1}\end{pmatrix},\quad A\in\mathrm{GL}_{g}({\mathcal{O}}_{k})\penalty 10000\ ,$
$n(B)$ $:=\begin{pmatrix}\mathrm{I}_{g}&B\\ 0&\mathrm{I}_{g}\end{pmatrix},\quad B\in\mathrm{Herm}_{g}({\mathcal{O}}_{k})\penalty 10000\ ,$
$w_{g}$ $:=\begin{pmatrix}0&\mathrm{-}I_{g}\\ \mathrm{I}_{g}&0\end{pmatrix}\penalty 10000\$.
@@
$\rho_{L,g}(m(A))\mathfrak{e}_{\underline{\nu}}$ $=\det(A)^{-p-q}\mathfrak{e}_{\underline{\nu}A^{-1}}\penalty 10000\ ,$
@@
for every element $T=\begin{pmatrix}A&B\\ C&D\\ \end{pmatrix}\in\mathrm{U}(g,g)(\mathbb{Z})\penalty 10000\ .$
2.2. Intertwining operators of Weil representations
Let $(L,h)$ be a Hermitian lattice over ${\mathcal{O}}_{k}$ as in the previous section and let ${\mathfrak{J}}\subset L$ be a maximal isotropic sublattice. Then $M={\mathfrak{J}}^{\bot}/{\mathfrak{J}}$ admits an induced Hermitian form, still denoted $h$, which is non-degenerate. The corresponding Weil representation will be denoted $\rho_{M,g}$.
Let $H$ be the image of ${\mathfrak{J}}_{k}\cap L^{\vee}$ in $L^{\vee}/L$ and let $H^{\bot}$ be its orthogonal with respect to the pairing on $L^{\vee}/L$. It is straightforward to check that the natural projection map $p:H^{\bot}\rightarrow M^{\vee}/M$ is surjective with kernel equal to $H$. Define the following maps:
$p_{L}^{M}:\,\mathbb{C}[(L^{\vee}/L)^{g}]$ $\rightarrow\mathbb{C}[(M^{\vee}/M)^{g}]$
(2.1) $\mathfrak{e}_{\underline{\nu}}$ $\mapsto\begin{cases}\mathfrak{e}_{p(\underline{\nu})}&\text{if $\underline{\nu}\in(H^{\bot})^{g}$,}\\ 0&\text{otherwise,}\end{cases}$
$p_{M}^{L}:\,\mathbb{C}[(M^{\vee}/M)^{g}]$ $\rightarrow\mathbb{C}[(L^{\vee}/L)^{g}]$
(2.2) $\mathfrak{e}_{\underline{\mu}}$ $\mapsto\sum_{\underline{\nu}\in H^{\bot},\,p(\nu)=\mu}\mathfrak{e}_{\nu}.$
The operators $p_{M}^{L}$ and $p_{L}^{M}$ intertwine the Weil representations on both sides, and hence we get a linear map between the corresponding spaces of vector-valued Hermitian modular forms.
@@ 2.3. Homogeneous polynomials and modularity
$P(UA)=|\det(A)|^{2}P(U),\quad\textrm{for all}\quad U\in\mathrm{M}_{n\times g}(\mathbb{C}),\,A\in\mathrm{M}_{g}(\mathbb{C})\penalty 10000\ .$
@@
$P(\lambda_{1},\ldots,\lambda_{g})=Q(\lambda_{1},\ldots,\lambda_{g},\overline{\lambda}_{1},\ldots,\overline{\lambda}_{g})\penalty 10000\ .$
@@
$\lambda_{j}\cdot\frac{\partial P}{\partial\lambda_{i}}=\delta_{i,j}P\quad,\quad\overline{\lambda}_{j}\cdot\frac{\partial P}{\partial\overline{\lambda}_{i}}=\delta_{i,j}P\penalty 10000\ .$
@@
where $Q(X_{1},\ldots,X_{g},Y_{1},\ldots,Y_{g})$ satisfies the condition in Equation˜2.1. We then apply [Roe21, Proposition 3.4]
@@
$\lambda_{\ell}=\begin{pmatrix}\lambda_{\ell}^{(1)}\\ \vdots\\ \lambda_{\ell}^{(n)}\end{pmatrix}\in\mathbb{C}^{n},$
@@
$\frac{\partial}{\partial\lambda_{\ell}}=\begin{pmatrix}\frac{\partial}{\partial\lambda_{\ell}^{(1)}}\\ \vdots\\ \frac{\partial}{\partial\lambda_{\ell}^{(n)}}\end{pmatrix}\penalty 10000\$.
@@
    For any $1\leq s,\ell\leq g$, we have: $\Delta_{s}P=\Delta_{\ell}P$ in ${\mathcal{F}}_{n,g-1}$ as functions of g1g-1 vectors.
@@
$P(\lambda_{1},\ldots,\lambda_{g})=Q(\lambda_{1},\ldots,\lambda_{g},\overline{\lambda}_{1},\ldots,\overline{\lambda}_{g})\penalty 10000\ .$
@@
$\Delta_{g}P(\lambda_{1},\ldots,\lambda_{g})=\sum_{s=1}^{n}\frac{\partial^{2}Q}{\partial X_{g}^{(s)}\partial Y_{g}^{(s)}}(\lambda_{1},\ldots,\lambda_{g},\overline{\lambda}_{1},\ldots,\overline{\lambda}_{g})\penalty 10000\ ,$
@@
$Q\left[(X_{1},\ldots,X_{g-1})\cdot A,X_{g},(Y_{1},\ldots,Y_{g-1})\cdot B,Y_{g}\right]\\ =Q\left[(X_{1},\ldots,X_{g})\cdot A_{0},(Y_{1},\ldots,Y_{g})\cdot B_{0}\right]\penalty 10000\ \\ =\det(A_{0})\det(B_{0})Q(X_{1},\ldots,X_{g},Y_{1},\ldots,Y_{g})\\ =\det(A)\det(B)Q(X_{1},\ldots,X_{g},Y_{1},\ldots,Y_{g})\penalty 10000\ .$
@@
For $(2)$, without loss of generality, we can assume that $s=g$ and $\ell=g-1$. Let $\sigma$ be the matrix of the transposition (g1,g)(g-1,g). Let $u_{1},\ldots,u_{g-1},\lambda_{g}\in\mathbb{C}^{n}$. The invariance property applied to $\sigma$ ensures that:
$P(u_{1},\ldots,u_{g-1},\lambda_{g})=P(u_{1},\ldots,\lambda_{g},u_{g-1})\penalty 10000\ .$
@@
$\Delta_{g}P(u_{1},\ldots,u_{g-1})=\Delta_{g-1}P(u_{1},\ldots,u_{g-1})\penalty 10000\ ,$
@@
Lemma˜2.6 above ensures that there exists a well-defined lowering operator for $1\leq g\leq n$:
$\Delta_{g}:{\mathcal{F}}_{n,g}\rightarrow{\mathcal{F}}_{n,g-1}\penalty 10000\ .$
@@
$\Delta:{\mathcal{F}}_{n,\bullet}\rightarrow{\mathcal{F}}_{n,\bullet}\penalty 10000\ ,$
@@
$m_{ij}=-\prescript{t}{}{\lambda_{i}}\cdot\frac{\partial m_{ii}}{\partial\lambda_{j}}=\sum_{s=1}^{n}{\lambda}_{i}^{(s)}\frac{\partial m_{ii}}{\partial\lambda_{j}^{(s)}}\penalty 10000\ .$
Using the polynomial $Q$ associated to $P$ as in Equation˜2.1 and applying the invariance property for permutation matrices, we can write the $m_{ij}$ as follows:
@@
where we introduce the following notation: for $g$ vectors $\mu_{1},\ldots,\mu_{g}\in\mathbb{C}^{n}$: $\mu]_{i}^{g}$ is the ordered set of g1g-1 vectors $(\mu_{1},\ldots,\widehat{\mu}_{i},\ldots\mu_{g})$ skipping $\mu_{i}$.
@@
$\lambda_{j}\cdot\frac{\partial m_{\ell k}}{\partial\lambda_{i}}=\lambda_{j}\cdot\frac{\partial}{\partial\lambda_{i}}(P(\lambda_{1},\ldots,\widehat{\lambda}_{\ell},\ldots,\lambda_{g}))\penalty 10000\ ,$
which 0 if $\ell=i$ (and therefore $k\neq j$). If $\ell\neq i$ and $k\neq j$, then we get 0 by Lemma˜2.5. Finally, if $\ell=j(=k)$, then we get $m_{j,i}=-m_{\ell,i}$.
Assume now that $k\neq\ell$, then $m_{\ell,k}=-\lambda_{\ell}\cdot\frac{\partial m_{\ell\ell}}{\partial\lambda_{k}}$. If $j\neq k$, then $\lambda_{j}\cdot\frac{\partial m_{\ell k}}{\partial\lambda_{i}}=0$, as $m_{\ell,k}$ either does not depend on $\lambda_{i}$, or we can apply Lemma˜2.5. Therefore, we may assume that $j=k$. By applying Equation˜2.1 to permutation matrices, we get:
@@
$\Lambda(P)(\lambda_{1},\ldots,\lambda_{g})=\mathrm{Tr}\left({{}^{t}{U}}\overline{U}M\right)\penalty 10000\ ,$
@@
It is enough to check that $\Lambda P$ satisfies the conditions of Lemma˜2.5. We can expand
$\Lambda(P)=\sum_{1\leq\ell,s\leq g}h(\lambda_{s},\lambda_{\ell})m_{\ell,s}\penalty 10000\ .$
@@
$\lambda_{j}\cdot\frac{\partial\Lambda(P)}{\partial\lambda_{i}}=\sum_{1\leq\ell,k\leq g}\lambda_{j}\cdot\frac{\partial}{\partial\lambda_{i}}\left(h(\lambda_{k},\lambda_{\ell})\right)m_{\ell,k}+\sum_{1\leq\ell,k\leq g}h(\lambda_{k},\lambda_{\ell})\lambda_{j}\cdot\frac{\partial m_{\ell,k}}{\partial\lambda_{i}}\penalty 10000\ .$
@@
$\lambda_{j}\cdot\frac{\partial}{\partial\lambda_{i}}(h(\lambda_{k},\lambda_{\ell}))=\delta_{ik}h(\lambda_{j},\lambda_{\ell})\penalty 10000\ .$
We distinguish two cases: assume first that $j=i$. Then by Lemma˜2.7, (1), we have:
$\lambda_{i}\cdot\frac{\partial m_{\ell,k}}{\partial\lambda_{i}}=(1-\delta_{ik})m_{\ell k}\penalty 10000\ .$
@@
$=\Lambda P\penalty 10000\ .$
We now treat the case $i\neq j$. By Lemma˜2.7 (2), we have: $\lambda_{j}\cdot\frac{\partial m_{\ell,k}}{\partial\lambda_{i}}=0$ if $k\neq j$ and otherwise it is equal to $m_{\ell,i}$. Therefore, we get:
@@
$=0\penalty 10000\ .$
We compute similarly $\overline{\lambda}_{i}\cdot\frac{\partial}{\partial\overline{\lambda_{j}}}\Lambda(P)=\delta_{ij}\Lambda P$. By Lemma˜2.5, we get the desired result. ∎
@@
$\Lambda,\Delta]=H,\quad[H,\Lambda]=2\Lambda,\quad[H,\Delta]=-2\Delta\penalty 10000\ .$
@@
(2.4) $\Lambda\circ\Delta_{g}(P)=\sum_{1\leq k,\ell\leq g}h(\lambda_{k},\lambda_{\ell})m_{\ell,k}(P_{g})\penalty 10000\ .$
@@
$+\sum_{1\leq k,\ell\leq g+1}\left[{\frac{{}^{t}\partial}{\partial\lambda_{g+1}}}h(\lambda_{k},\lambda_{\ell})\frac{\partial}{\partial\overline{\lambda}_{g+1}}m_{\ell,k}(P)+{\frac{{}^{t}\partial}{\partial\overline{\lambda}_{g+1}}}h(\lambda_{k},\lambda_{\ell})\frac{\partial}{\partial\lambda_{g+1}}m_{\ell,k}(P)\right]\penalty 10000\ ,$
@@
    $m_{g+1,g+1}(P)=P$, $\Delta_{g+1}(h(\lambda_{k},\lambda_{\ell}))=n$, if $k=\ell=g+1$ and 0 otherwise.
@@
    $\Delta_{g+1}(m_{g+1,k})=\Delta_{g+1}(m_{\ell,g+1})=0$ for all $1\leq\ell,k\leq g+1$, as $m_{g+1,k}$ and $m_{\ell,g+1}$ depend linearly on $\lambda_{g+1}$ and $\overline{\lambda}_{g+1}$ respectively.